OpenStudy (anonymous):

PRECALCULUS HELP ! domain & range of rational function 2 f(x) = --------- x^2-2x-3 i got domain = all real x except -1 and 3. i dont know how to find range please help

OpenStudy (anonymous):

it has a horizontal asymptote at y=0 so y > 0 is part of range

OpenStudy (anonymous):

but i dont know if thats correct way

OpenStudy (anonymous):

To find the horizontal asymptotes, you look at: \[\frac{\text{Degree of Numerator}}{\text{Degree of Denominator}}=\frac{0}{2}=0\] So there is a horizontal asymptote at \(y=0\). Thus this is not included in the range, and the range is: \[y\in(-\infty,0)\cup (0,\infty) \] Does that make sense?

OpenStudy (anonymous):

yes i thought the same initially... but it has vertical asymptotes also, when i graphed it in geogebra i get confuse because in the middle it is stopping at -1/2

OpenStudy (anonymous):

i have been struggling to figure out how to get -1/2 as upper limit in range y > 0 or, y <= -1/2 http://www.wolframalpha.com/input/?i=2%2F%28x%5E2-2x-3+%29

OpenStudy (anonymous):

Sorry, it's early I seem to have made a mistake. This document may help you out: http://cims.nyu.edu/~kiryl/Precalculus/Section_3.7-Rational%20Functions/Rational%20Functions.pdf I'll have a look and see if I can explain it at all.

OpenStudy (anonymous):

its okay... im going through the pdf thnks for the help xD

Directrix (directrix):

@sara12345 Can you show that -1/2 is the *only* relative maximum of the function? We know the function cannot have a value of 0. That will leave a gap in the Real range between [-1/2 , 0). Just thinking.

OpenStudy (anonymous):

i got y-intercept at -2/3, wil it be any use ?

Directrix (directrix):

I don't think so unless you can show -2/3 is a relative maximum. Can you get -1/2 as the relative maximum?

OpenStudy (anonymous):

sorry m not sure what is relative maximum

OpenStudy (anonymous):

to find the range put x=y in the equation you and then solve

OpenStudy (anonymous):

2 ------- y^2-2y-3

Directrix (directrix):

@sara12345 See attached file.

OpenStudy (anonymous):

oh yes that maximum only geogebra shows at -1/2, how to find it ?

OpenStudy (anonymous):